OpenStudy (anonymous):

Evaluate: 2^3/(2^0 - 2^-1)

OpenStudy (anonymous):

8 / (1 - 1/2) = 8 / 1/2 = 16

OpenStudy (anonymous):

okay thanks :) but how do you know that 2^-1 is = 1/2

OpenStudy (anonymous):

a negative exponent means 1 / +ve so 3^-1 = 1/3^1 = 1/3 3^-2 = 1/3^2

OpenStudy (anonymous):

Okay :) I get it now. THanks :)

OpenStudy (anonymous):

Also... do you know if I want to make the exponent -2^-1 apositive exponent and I flipped it to the top of the equation, would the two become positive????